### v1

# Post-glitch Recovery and the Neutron Star Structure: The Vela Pulsar

Abstract
Pulsar glitches, sudden changes in a neutron star’s rotation, offer a unique window into the extreme physics governing these celestial bodies. The post-glitch recovery phase, characterized by a gradual recovery, reveals insights into the superfluid dynamics within the star. The post-glitch recovery phase probes the coupling mechanisms between different stellar components, sheds light on the properties of neutron star matter at supra-nuclear densities and even allows us to predict the time to the next glitch. In this work, we present a detailed analysis of the Vela pulsar’s rotational behavior using around 100 months of observational data spanning from September 2016 to January 2025, during which four glitches were identified. Here, we demonstrate the post-glitch recovery of these glitches within the framework of the vortex creep model, providing new insights into the Vela pulsar’s internal structure. Notably, we present the first-ever investigation of vortex residuals – the discrepancy between observed values and those predicted by the vortex creep model – through the lens of the vortex bending model, marking a significant step forward in understanding neutron star physics through pulsar glitches.
I Introduction
Pulsars are highly magnetized rotating dense and compact stars observed to have periodic radiation of pulses with the periodicity tied to its remarkably stable rotation. The slow-down in the rotation of a pulsar is influenced by two types of instabilities: glitches and timing noise (D’alessandro, 1996; Haskell and Melatos, 2015; Zhou et al., 2022). Glitches are abrupt changes in the rotational frequency and spin-down rate of a pulsar, while timing noise is a stochastic wander in the spin parameters of pulsars. The occurrence of glitches, as well as their slow relaxations ranging from a few days to several years, reveal the presence of neutron superfluidity inside the inner crust of the neutron star. The most successful mechanism explaining glitches and their recoveries is based on the superfluid vortex pinning and creep model (Anderson and Itoh, 1975; Alpar et al., 1984a, b). This model suggests that neutron superfluidity in the interior of the star leads to a differential rotation, storing a significant amount of angular momentum. When this angular momentum is released, the crust spins up, resulting in a glitch. Pulsar glitches thus provide valuable insights into the superfluid dynamics within the star (Zhou et al., 2022).
The post-glitch recovery is the process in which the pulsar gradually relaxes back towards its original pre-glitch rotational state. According to the vortex creep model, this process is driven by the repinning of superfluid vortices to the crustal nuclei and the re-establishment of the equilibrium lag between the rotational rates of the superfluid interior and the crust. Hence, the vortex dynamics play a crucial role in determining the recovery behavior. The post-glitch relaxation typically involves the contribution of angular momentum transfer from the neutron star interior to the surface, seen as an exponential recovery followed by a linear decay in the spin-down rate. Occasionally, some inward movement of vortices may also play a role (Akbal et al., 2015) where crustquakes initiate glitches. The linear type of relaxation in time originates from the tangling or bending of the vortices in the non-linear creep regions (Gügercinoğlu et al., 2023). The post-glitch recovery can be utilized to probe the internal dynamics and structure of neutron stars (Gügercinoğlu et al., 2022).
PSR J0835–4510, B0833–45 or the Vela pulsar, has shown 26 glitches, most of which are large glitches and detected after a regular, quasi-periodic inter-glitch time. The Vela pulsar is one of the most studied glitching pulsars. The first-ever glitch detection was recorded in the Vela pulsar (Radhakrishnan and Manchester, 1969; Reichley and Downs, 1969). Moreover, the first-ever glitch detection caught in the act was also realized for the Vela pulsar (Palfreyman et al., 2018). The Vela pulsar is titled the most popular or most studied glitching pulsar among other glitching pulsars for various reasons: it is a bright source, young, and displays glitches at regular intervals. The recoveries of glitches also demonstrate strikingly similar behavior. A measure of a pulsar’s slow down over time is described by the braking index (). It is defined by a power-law relation between the pulsar’s spin-down rate and its rotational angular frequency (), given by . The braking index provides insights into the pulsar’s rotational evolution, magnetic fields, and emission mechanisms.
Here, we present the observational behavior of the Vela pulsar from September 2016 to January 2025; during this period, four large glitches were observed in our monitoring campaign (Grover et al., 2024a; Basu et al., 2020). Theoretical inferences from the post-glitch recovery phase are also presented. We have utilized the recovery behavior of four glitches in the Vela pulsar to estimate the following parameters: (a) the fractional moments of inertia (FMI) of various layers participating in a glitch, (b) the re-coupling timescale of the crustal superfluid, (c) theoretical prediction for the time to the next glitch, (d) braking index of the Vela pulsar. The paper’s outline is as follows: we describe the pulsar observations, the data reduction and the analysis processes in section II. A brief description of the vortex creep and bending models is given in Section III. In Section IV, we discuss the results. The conclusions and future work are given in Section V.
II Observations and Analysis
II.1 Sample and Observational Facilities
We use two large radio telescopes to monitor pulsars: the upgraded Giant Metrewave Radio Telescope (uGMRT)(Gupta et al., 2017) and the Ooty Radio Telescope (ORT) (Swarup et al., 1971). Observations of PSR B083345 were part of a regular glitch monitoring program, where we regularly monitor 24 pulsars (i) to detect glitches, (ii) model timing noise, and (iii) model post-glitch recovery. This work presents a comprehensive analysis of the Vela pulsar’s rotational evolution, spanning from September 2016 to January 2025, as part of our ongoing monitoring campaign (Grover et al., 2024a; Basu et al., 2020). Our study primarily utilizes observations from the ORT, supplemented by uGMRT and archival ATNF data to fill temporal gaps and facilitate specialized investigations.
The ORT (Swarup et al., 1971) is a 530m 30m parabolic cylindrical reflector. It is located on a hill with the same slope as the geographic latitude (11°), allowing it to track sources for about 8 hours and covers a declination range of -60° to +60°. Pulsar observations at the ORT are carried out at the central observation frequency of 326.5 MHz using the Pulsar Ooty Radio Telescope New Digital Efficient Receiver, PONDER (Naidu et al., 2015), which produces real-time coherent dedispersed time series data. The typical cadence of the observations at the ORT is 1–14 days.
The GMRT is an interferometer with 30 parabolic antennas, with a diameter of 45 meters each (Swarup et al., 1991; Ananthakrishnan, 1995). The uGMRT (Gupta et al., 2017) covers a declination range of -53°to +90°and supports observations from 150 MHz to 1460 MHz in four bands. The Vela pulsar is observed in uGMRT Band 3 (250–500 MHz) and Band 4 (550–950 MHz). The central square antennas with a subset of nearest arm antennas were used in a phased array synthesizing a large single dish for observations. The outer arm antennas were not included in the phased array due to significant drift in the phase of the voltage outputs, caused by ionospheric variations, in these antennas. A digital filterbank, implemented in the GMRT correlator, was used to record either 1024 or 2048 channel across the bandpass to obtain frequency resolved data with a typical sampling time of 655 s. The cadence of the observations is 15-30 days. We utilize the frequency-resolved uGMRT data primarily to examine the impact of dispersion measure (DM) offsets on the spin-down rate. Additionally, uGMRT data filled a temporal gap around epoch 60300, providing crucial estimates not covered by ORT due to maintenance downtime.
The Commonwealth Scientific and Industrial Research Organisation (CSIRO) Data Access Portal is used to search and access the Parkes pulsar archival observations111https://data.csiro.au/domain/atnf (Hobbs et al., 2011) of the Vela pulsar. These data were mainly used to fill the temporal gaps in glitch number 2, or MJD 58515 glitch, caused by significant observational gaps at ORT. We directly downloaded and utilized the reduced files (or PSRFITS (Hotan et al., 2004) files) and performed subsequent analysis, as described in the next subsection.
II.2 Data Reduction and Analysis
We utilized the DSPSR software (van Straten and Bailes, 2011) and the PINTA pipeline (Susobhanan et al., 2021) to reduce the raw data of the Vela pulsar observed at the ORT and the uGMRT, respectively, into the PSRFITS files with the ephemeris created in our initial observations. The initial analysis, including the generation of a noise-free template and Time of Arrivals (ToAs), was performed using PSRCHIVE (Hotan et al., 2004; van Straten et al., 2012). The ToAs were obtained through the cross-correlation of the template profile with all other observed profiles using the ‘pat’ command, utilizing the Fourier Phase Gradient (PGS) method (Taylor, 1992). The precision timing analysis of the pulsar was done with the help of the pulsar timing software TEMPO2 (Hobbs et al., 2006; Edwards et al., 2006). The DM offsets were estimated using the DMcalc (Krishnakumar et al., 2021) script from the frequency-resolved data from uGMRT.
The rotational evolution of a pulsar during a glitch can be expressed as a Taylor series expansion,
where is the spin frequency, and and are the first and second derivatives of the frequency, respectively, at reference epoch . The spin evolution of the Vela pulsar from September 2016 to January 2025 obtained after precision timing is given in Fig. 1. Each measurement is made with the local timing of several epochs, typically within a 50-200 day data window, contingent on data availability. Exceptions include the first and last few points adjacent to glitches. The resulting measurements are reflected in the precision of error bars, which are smaller than the symbol used in the figure. By choosing multi-epoch windows, we achieve reliable rotational estimates with minimized error bars, providing a robust foundation for studying vortex residuals. This approach effectively averages out temporal effects caused by physical processes, such as ionized interstellar medium scattering-induced broadening, thereby validating the accuracy of measurements aligned with the star’s rotation and internal dynamics.
For each recovery, we utilized Bayesian statistics for model comparison and parameter estimation. Bayesian statistics uses Bayes’ theorem to enhance the probability of the hypothesis based on new evidence or data. For model comparison, the standard Jeffreys scale is utilized, which suggests the preferred model based on the strength of evidence (Trotta, 2008). Bayesian inference is also used to predict the epoch of the next glitch using data from previous glitches.
The priors are selected based on two criteria: initial estimates from the weighted least-square fit and reasonable ranges for each parameter using literature knowledge. These reasonable ranges can be fixed for some parameters, like the fractional moment of inertia (typically positive and below 1) and may vary for others like the recoupling timescale, usually depends whether it is a 3 or 2 or 1 exponential component model. For the linear relaxation region, the prior for recoupling timescale is 10-300 days, and the offset time is 300 to 1200 or more for some recoveries with the initial guess as given by the weighted least-square technique. All the priors were considered as uniform priors. Since all the recoveries are unique hence this flexible approach to setting priors is necessary. The Gaussian likelihood is considered for the analysis. The MCMC Hammer: emcee (Foreman-Mackey et al., 2013) ensemble sampler is used for both parameter estimation and model comparison.
The post-glitch recoveries are studied within the framework of the vortex creep model and the vortex bending model (Gügercinoğlu et al., 2023). We first model the frequency derivative measurements from the data using the vortex creep model. Then, we estimate the residuals obtained after subtracting the predicted value by the vortex creep model from the measured values. These residuals were investigated within the framework of the vortex bending model. The generalized Lomb-Scargle periodogram technique (Lomb, 1976; Scargle, 1982; VanderPlas, 2018) was employed to detect periodicities in the residuals. Peaks with false alarm probabilities (FAP) below 5% or Z-score above 3 were considered significant for vortex bending studies. FAP estimation was performed using the Astropy library (Astropy Collaboration et al., 2013), which implements the Baluev (2008) approach. Alternative methods, including approximations based on effective independent frequencies and bootstrap resampling, yielded similar results. The Z-score significance is obtained from the tail of the one-sided Gaussian distribution corresponding to the observed FAP (Cowan et al., 2011; Ganguly and Desai, 2017). The formalism of the vortex creep and bending models is described in the next section.
III The Vortex Creep and Bending Models
The Vortex Creep model describes the occurrence of glitches and the post-glitch recovery as a byproduct of the superfluid present inside the star. The important features of vortex creep for this paper are now described.
In the neutron star crust, vortex lines coexist with lattice nuclei, and it is energetically favorable for the vortices to pin to the crystal lattice nuclei. The distribution of vortices is non-uniform due to crustal inhomogeneities such as dislocations, cracks, and impurities formed after quakes. This results in formation of structures such that regions of high vortex density are surrounded by areas with vortex depletion regions collectively referred to as vortex traps (Cheng et al., 1988). Glitches may arise due to collective vortex unpinning events induced in these vortex traps when the angular velocity lag between the superfluid and the crustal normal matter exceeds a threshold value.
After a glitch, the steady-state lag decreases, causing the cessation of creep in certain superfluid regions. The post-glitch recovery process involves the restoration of the creep until the steady-state conditions are re-established by the external braking torque. The vortex creep model describes the formation, recovery and other phenomena related to glitches as follows (Alpar et al., 1984a; Gügercinoğlu and Alpar, 2020; Gügercinoğlu et al., 2022):
A crustquake initiates the formation of vortex traps and triggers a vortex unpinning avalanche. The unpinned vortices transfer their excess angular momentum to the crust, causing it to spin up, which is observed as a glitch.
Superfluid regions in which unpinning occurs decouple from the external decelerating torque, resulting in a step increase in the spin-down rate.
Soon after the large-scale vortex unpinning avalanche, the excess angular momentum of the crustal superfluid is first transferred to the normal matter nuclei and electrons in the crust. This transfer causes a peak in the crustal rotation rate following a glitch.
This change in the crustal rotational state is then communicated to the core superfluid via electrons scattering off magnetized vortex lines (Alpar et al., 1984b). Consequently, the initial peak in the observed crustal rotation rate relaxes on the crust-core coupling time scale, eventually reaching the post-glitch equilibrium rotation rate.
A peak in the rotation rate of the normal component will occur if the coupling between the crustal superfluid and the normal matter is stronger than the coupling between the crustal superfluid and the core superfluid (Pizzochero et al., 2020).
The slow post-glitch evolution is the response of the crustal superfluid to the glitch-induced changes. On longer time scales, the core superfluid reaches equilibrium with the normal matter in the crust, and they behave as a single component. Consequently, the only component decoupled from the braking torque is the crustal superfluid.
The crust breaking can produce transient magnetospheric signatures that coincide with the glitch, such as changes in pulse shape, polarization, and emission. These changes are linked to the plastic motion of the magnetic field lines anchored to the fractured crustal platelets.
Damped sinusoidal-like oscillations in the spin-down rate can appear due to the bending of the vortex creep lines in the linear creep relaxation region, which is mainly generated by the shocks from the respective glitch and can also have an impact from the previous glitches (Gügercinoğlu et al., 2023).
The post-glitch spin-down rate behavior mirrors the responses of the vortex creep regions in both linear and non-linear regimes to a glitch, with additional involvement of induced inward vortex motion in delayed spin-up events. In certain regions of the superfluid in the inner crust and the outer core, vortex creep exhibits a linear response to glitch-induced changes, leading to exponential relaxation (Alpar et al., 1989; Flanagan, 1990; Gügercinoğlu, 2017) as:
where , is the spin frequency, is the spin frequency just before the glitch, is the FMI of the of ith layer that participated in a glitch, and is the exponential recovery timescale.
The exponential decay time scale in the linear creep region of the crustal superfluid relates with the temperature , pinning energy between the vortex line and nucleus , the critical lag for unpinning , the superfluid angular velocity , the Boltzmann constant , and the distance of vortex lines from the rotation axis as (Alpar et al., 1989):
where the microscopic vortex velocity around nuclei, m s-1 throughout the crust (Gügercinoğlu and Alpar, 2016).
The post-glitch response in the non-linear regime for a single crustal superfluid region is (Alpar et al., 1984a):
where is the FMI of the single crustal superfluid region in the non-linear regime, is the pre-glitch spin-down rate, is the recoupling timescale for the single crustal superfluid region in the non-linear regime and the offset time is .
The overall linear relaxation can be obtained by integrating the contributions of several neighboring single non-linear regime regions, i.e., equation 4 with an assumption of linearly decreasing superfluid angular velocity during a glitch (Alpar et al., 1996) and is given as:
where is the FMI of the non-linear crustal superfluid region, is the offset time, and is the non-linear creep relaxation time.
The recoupling timescale in the non-linear creep region of the crustal superfluid relates with the temperature , pinning energy between the vortex line and nucleus , the critical lag for unpinning , the superfluid rotational velocity rate , and the Boltzmann constant as (Alpar et al., 1984a):
The non-linear creep response gives an estimate for the time to the next glitch. The glitch-induced changes in the spin-down rate fully relax after a waiting time (Gügercinoğlu et al., 2022),
where is the observed glitch amplitude and represent the decrement in the superfluid rotational velocity. The post-glitch recovery completes once the relaxation of the non-linear creep regions is finished. This gives us a theoretical prediction for the time to the next glitch in terms of the vortex creep model, which lies between the range to .
An external effect like a crustquake can add an inward vortex motion. This increases the superfluid rotation rate as the vortices pinned to the broken plate are induced to move towards the rotation axis along with the plate. When vortices move inward, the superfluid rotates faster. The lag increases from its steady-state value, making the creep more efficient than in steady state and resulting in an enhanced vortex creep current in the radially outward direction (Gügercinoğlu and Alpar, 2019). The post-glitch response for inward moving vortices causing delayed spin-up observed from some glitches is given as (Akbal et al., 2015):
where is the FMI of the non-linear creep superfluid region in which inward vortex motion took place, and and are corresponding recoupling timescale and offset time associated with inward vortex motion.
A typical post-glitch recovery involves an exponential and a linear relaxation described by equations 2 and 5, respectively. A post-glitch recovery where delayed spin-ups have been observed requires an additional non-linear regime response term provided by equation 9. One or more of these is relevant to the post-glitch behaviour of the Vela pulsar investigated in this paper. This is tested by using different alternative models as described below:
Model 1: Only exponential recovery given by equation 2.
Model 1a: Exponential recovery with only 1 exponential component.
Model 1b: Exponential recovery with 2 exponential components.
Model 1c: Exponential recovery with 3 exponential components.
Model 2: Only linear relaxation described by equation 5.
Model 3: Exponential and linear recoveries, the sum of equations 2 and 5.
Model 3a: The sum of linear recovery and exponential recovery with only 1 exponential component.
Model 3b: The sum of linear recovery and exponential recovery with 2 exponential components.
Model 3c: The sum of linear recovery and exponential recovery with 3 exponential components.
The evidence for these different alternatives was evaluated using Bayesian analysis of our post-glitch recovery measurements of the Vela pulsar to select the suitable model and estimate its parameters.
Apart from long-term post-glitch recovery, we see evidence for sinusoidal-like spin-down rate oscillations over and above the recovery predicted by the vortex creep model as described in Section IV. Such oscillations can be explained by vortex bending (Gügercinoğlu et al., 2023) and appear to be a common feature following glitches, as also observed from PSRs J1522–5735 (Zhou et al., 2024), J0742–2822 (Zubieta et al., 2025a), J1509–5850 and J1718–3825 (Liu et al., 2025). To understand this better, we used three alternative hypotheses as described below
Hypothesis I – An undamped model consisting of only sinusoidal term, .
Hypothesis II – Damped model consisting of a sinusoidal term and a damping factor, .
Hypothesis III – An oscillatory model as described in Gügercinoğlu et al. (2023) consisting of .
The evidence for these hypotheses was also evaluated using Bayesian analysis to select the most suitable model and calculate its parameters. The bending of vortices remains a relatively unexplored domain, yet it can significantly impact simulations of neutron star glitches. Incorporating oscillation parameters associated with vortex bending, as reported in the next section, may provide a more comprehensive understanding of glitch behavior and boost the hydrodynamic simulation of the superfluid vortices (Howitt et al., 2020; Anantharaman and Bhattacharya, 2025).
IV Results and Discussions
Our pulsar monitoring program with the ORT and the uGMRT started in 2015. The previous results were reported in Basu et al. (2020) and Grover et al. (2024a). In this work, we present analysis of a new Vela glitch at MJD 60429.9. Further, we demonstrate the spin-down rate evolution for the Vela pulsar since the beginning of the monitoring program. Our primary focus is to provide the theoretical interpretation of glitches using the post-glitch recovery phase and to obtain measurements for the re-coupling timescales and fractional moment of inertia for various layers that participate in a glitch. We have estimated the theoretical time to the next glitch through the vortex creep model and used Bayesian statistics to predict the time for the next glitch. All the predictions match well with the observed inter-glitch times.
IV.1 Glitches
Here, we summarize the last three glitches in the Vela pulsar: MJD 57734.4, 58517, and 59417.6, detected in our monitoring program (Grover et al., 2024a; Basu et al., 2020). The recoveries for each of these glitches are studied within the framework of the vortex creep model to investigate the structure of the Vela pulsar. Furthermore, we report the detection and analysis of the recent large glitch in the Vela pulsar in 2024. The characteristics and parameters of glitches are listed in Table 1.
| Number | Epoch (MJD) | Pre-glitch (Hz) | Pre-glitch (Hz s-1) | () | Reference | |
| G1 | 57734.4(2) | 11.186433252(6) | -1.556383(8) | 1433.2(9) | 5.595(9) | Basu et al. (2020) |
| G2 | 58517(7) | 11.185397242(3) | –1.555659(4) | 2471(6) | 6(2) | Grover et al. (2024a) |
| G3 | 59417.6(1) | 11.184208478(5) | –1.5550(8) | 1235(5) | 8.0(7) | Grover et al. (2024a) |
| G4 | 60429.9(1) | 11.182859362(3) | –1.5530(8) | 2396(2) | 23(1) | This work |
We present the parameters of the Vela 2024 glitch, which occurred around MJD 60429.9. This glitch has been reported by several groups (Campbell-Wilson et al., 2025; Zubieta et al., 2025b). Here, we present updates in our preliminary results reported in Grover et al. (2024b). The glitch is estimated to occur at MJD 60429.9(1), with the fractional increase in the spin frequency, and its time derivative is calculated as and , respectively. The results of glitch epoch and spin parameters agree with the reported estimates of other monitoring programs. A more precise measurement of glitch epoch as 60429.86962(4) is also presented in Palfreyman (2024) with higher cadence observations. The timing residuals, the evolution of the spin frequency and the spin-down rate are given in Fig. 2 and are being presented for the first time for this new glitch. The post-glitch recovery for this glitch is also investigated within the framework of the vortex creep model in the next subsection.
IV.2 Post-Glitch Recoveries
In this section, we present the post-glitch recovery behavior of the Vela pulsar. The spin-down rate has been analyzed from MJD 57651 to 60700. During this period, the Vela pulsar has displayed four glitches. The post-glitch recovery of the Vela pulsar is regular, i.e., it follows similar behavior. The recovery for this pulsar usually consists of exponential and linear relaxations well described by Model 3. We have investigated all post-glitch recoveries for this pulsar within the framework of the vortex creep model for the glitches listed in Table 1, and the resulting fit parameters obtained using the vortex creep model are listed in Table 2. Now, we concisely describe the post-glitch relaxation(s) for individual cases.
| Epoch | |||||||||||
| (MJD) | () | (days) | () | (days) | () | (days) | () | (days) | (days) | (days) | (days) |
| 57734.4(2) | — | — | 783 | [787, 1083] | |||||||
| 58517(7) | 902 | [752, 1033] | |||||||||
| 59417.6(1) | 1012 | [855, 1123] | |||||||||
| 60429.9(1) | — | [820, 1076] |
IV.2.1 Glitch No. 1 – MJD 57734.4(2)
G1, or glitch at MJD 57734.4(2) (Palfreyman et al., 2018; Basu et al., 2020), was detected with a fractional change in the rotational frequency of 1433(1), and a fractional change in the spin-down rate as 5.59(1). We present the post-glitch relaxation modeling for this glitch in Fig. 3. For this recovery, the Bayesian model comparison prefers the presence of two exponentials (Model 3b) over all other models. The estimated parameters are: the fractional moment of inertia for the second and third exponential components are and respectively, with the decay timescales of and days, respectively. The fractional moment of inertia for the non-linear creep region is with a recoupling timescale of 61 days. The offset time is measured as days, close to the observed inter-glitch time of 783 days. The observed time is consistent with the vortex creep model’s predicted range and closely matches the Bayesian prediction, which estimates a 68% credible interval of 787 to 1083 days. The fit parameters are listed in Table 2.
The bottom part of Fig. 3 shows the residuals obtained after subtracting the vortex creep model predictions from the observed data and is in units of Hz s-1. These vortex residuals, spanning from 200 days onward, were utilized for vortex bending studies, excluding the initial 200 days due to the significant impact of exponential recovery. The generalized Lomb-Scargle periodogram analysis, Fig. 4, revealed a prominent peak corresponding to a period of 314.1(2) days, with a false alarm probability (FAP) of (Z-score: 4.9). A secondary peak at approximately 166 days had a FAP of 0.194, and other subsequent peaks had FAP values of 1, indicating low significance. Fixing the period at 314.1(2) days, we investigated the vortex bending hypothesis and found that the damping model, comprising a sinusoidal term and a damping factor (Hypothesis II), was strongly favored over Hypotheses I and III. Additionally, we examined whether the observed peak resulted from shocks caused by the current glitch or the previous one, concluding that the oscillations decisively originated from the current glitch. The comparison of vortex residuals and vortex bending model is shown in Fig. 5 and the estimated parameters for the damped model (Hypothesis II) are: Amplitude = , Phase = , and days. Note that vortex residuals are in scale of Hz s-1, so the effective amplitude is .
The vortex residuals fit with the vortex bending model well for around 200-400 days post-glitch duration but deviate thereafter. Including a second peak improves the fit, but the generalized Lomb-Scargle periodogram indicates a FAP of 19.4% for this peak using the Baluev method (with a minimum FAP of 8.9% obtained through approximations based on effective independent frequencies). Given the relatively high FAP, we exclude the second component from our analysis. The other peaks with a high FAP are likely to originate from preceding glitches, and their effect on the present glitch recovery is insignificant. The post-glitch recovery modeling for this pulsar is also presented by Gügercinoğlu et al. (2022) using the Fermi-LAT observations. However, they only detected one exponential component. We are reporting 2 exponential recoveries made possible due to much higher cadence observations allowing us to investigate the vortex residuals in the framework of the vortex bending model for the first time.
IV.2.2 Glitch No. 2 – MJD 58517(7)
The glitch at MJD 58517(7), G2 in the Vela pulsar, (Grover et al., 2024a; Lower et al., 2020; Gügercinoğlu et al., 2022), was observed with a fractional change in the rotational frequency of 2471(6), and a fractional change in the spin-down rate as 6(2). Our analysis focuses on modeling the post-glitch relaxation behavior, given in Fig. 6. Observations weren’t conducted during the exponential relaxation phase of this glitch, and we detected only non-linear creep recovery. To proceed with our post-glitch analysis, we supplemented our dataset with publicly available ATNF pulsar data to fill temporal gaps. Additionally, we utilized estimates derived from analysis using data closer to the glitch from Lower et al. (2020), specifically the glitch epoch (58515.6) and the fractional change in rotational frequency and it’s derivative (2501 and 8.7, respectively), for our recovery studies.
The Bayesian model comparison reveals that the three exponential model (Model 3c) best describes this recovery. The measured values for the recovery parameters are as follows: the fractional moment of inertia for the first, second and third exponential components are and and respectively, with the decay timescales of , and days respectively. The fractional moment of inertia for the non-linear creep region is with a recoupling timescale of days. The offset time is estimated as days. The observed inter-glitch time of 902 days is consistent with both the predicted range of the vortex creep model and Bayesian statistics. The median value predicted by Bayesian statistics is 893 days, very close to the observed time. These parameters are also summarized in Table 2.
The first exponential recovery timescale is approximately 7 days, roughly 3 or more times larger than the corresponding component in the next two glitches. Given this disparity, we investigated the possibility of an additional exponential with a shorter timescale (0.1–2 days) but found no converging results, suggesting that the 7-day timescale is a distinct characteristic of the initial recovery phase. The third exponential component’s precision is biased due to poor constraint, affecting the measurements of vortex residuals. The third exponential recovery’s impact on vortex bending is its timescale of approximately 600 days; hence, we did not study vortex residuals for this glitch using the vortex bending model. Our post-glitch recovery modeling results differ from those of Gügercinoğlu et al. (2022), who reported a single exponential component using Fermi-LAT observations. Our analysis benefits from a higher cadence and reveals two well-constrained distinct exponential recoveries.
IV.2.3 Glitch No. 3 – MJD 59417.6(1)
G3, or the glitch at MJD 59417.6(1) (Grover et al., 2024a; Zubieta et al., 2023) was observed to have a fractional spin up in frequency and a fractional spin up in frequency-time derivative . We present the post-glitch relaxation behavior for this glitch in Fig. 7. In addition to the ORT, we leverage observed data from the uGMRT as well, with two motivations: (i) to confirm that the residuals stem from the neutron star’s interior, rather than the ionized interstellar medium by utilizing DM offsets in the uGMRT analysis and (ii) to fill the temporal gap between 850 to 1000 days subsequent to the glitch date.
High-cadence ORT observations of this glitch enable us to model three exponential components present in the recovery; Model 3c emerged as the top choice in Bayesian model comparisons, outperforming all other models. The calculated values for the parameters are as follows: the fractional moment of inertia for the first, second and third exponential components are and and respectively, with the decay timescales of , and days, respectively. The fractional moment of inertia for the non-linear creep region is with a recoupling timescale of days. The offset time is estimated as days, which deviates significantly from the observed inter-glitch time primary due to the large value of (equivalent to ). However, the vortex creep model predicted time for the next glitch lies between to , which aligns well with the observed time. The median Bayesian prediction of 989 days is near the observed interval of 1012 days, falling well within the 68% credible interval. These measured parameters are listed in Table 2.
The relaxation behavior of this glitch has never been investigated before. The lower panel of Fig. 7 shows the vortex residuals, obtained after subtracting the vortex creep model predictions from the observed data, in scale of Hz s-1. We analyzed residuals between 200 and 700 days for vortex bending studies, excluding the initial 200 days due to dominant exponential recovery effects and data beyond 700 days, which warrants deep investigation.
The Lomb-Scargle periodogram analysis of vortex residuals, depicted in Fig. 8, exhibits multiple peaks. However, only the first two peaks, corresponding to periods of approximately 344(6) and 153(3) days, demonstrate statistical significance. The first peak yields a FAP of 4.6 10-13 and a Z-score of 7.1, while the latter peak has a FAP of 2.46 10-5 and a Z-score of 4.1. The third peak, with a period of around 81 days, has a FAP of 0.88 and is not significant, nor other peaks all of which have a FAP of 1. These two peaks are used to investigate the vortex residuals within the framework of the vortex bending hypothesis, presented in Fig. 9. We observed that Hypothesis II was strongly favored over Hypotheses I and III. Additionally, we examined whether the observed peak resulted from shocks caused by the current glitch or the previous one, concluding that the oscillations decisively originated from the current glitch. The estimated parameters for the damped model (Hypothesis II) are: (i) Amplitudes, A1 = , and A2 =, (ii) Phase, and , (iii) and days. Note that vortex residuals are in scale of Hz s-1, so the effective amplitudes are , and , respectively.
IV.2.4 Glitch No. 4 – MJD 60429.9(1)
We present the post-glitch relaxation, presented for the first time, for the recent glitch at MJD 60429.9(2), G4 (Grover et al., 2024b; Campbell-Wilson et al., 2025; Zubieta et al., 2025b). The glitch plot is given in Fig. 2, and the comparison of the observed data with the vortex creep model fitting is presented in Fig. 10.
The Bayesian model comparisons selected the Model 3c as the best model over all other models. The calculated values for the parameters are as follows: the fractional moment of inertia for the first, second and third exponential components are and and respectively, with the decay timescales of , and days, respectively. The fractional moment of inertia for the non-linear creep region is with a recoupling timescale of days. The time for the next glitch is predicted as days. These parameters are listed in Table 2. The parameters for the linear creep regions (or exponential recovery) are well-constrained. However, those dependent on the non-linear creep region will likely be refined in the future as more data becomes available. The Bayesian analysis forecasts a median time for the next glitch at MJD 61377.7, with a 68% credible interval spanning from MJD 61249 to 61506.
IV.3 Correlations and Braking Index
To date, the Vela pulsar has exhibited 26 glitches, with 21 of them being with large magnitudes exceeding 10-8. Fig 11 presents these 21 events, revealing a pattern where glitch magnitudes exhibit an alternating increasing and decreasing trend: if glitch A has a lower magnitude, the successive glitch is observed with a higher magnitude, and vice versa. This trend is particularly evident in the last 10 glitches (or after MJD 50000).
We also utilized the data of these 21 events to investigate potential correlations between the glitch amplitude and the inter-glitch time. Most correlations were not statistically significant except a correlation between glitch amplitude and subsequent inter-glitch time, shown in the Fig 12. Specifically, the amplitude of a glitch (e.g., glitch A) showed a positive correlation with the inter-glitch time following it (i.e., the time difference between glitch A+1 and glitch A). The correlation coefficients were: Pearson’s = 0.6582 (P–value = 0.0022), Spearman’s = 0.6266 (P–value = 0.0041), and Kendall’s = 0.4692 (P–value = 0.0051), indicating moderate to strong positive correlation. Such a tight correlation has only been found for PSR J0537–6910 (Ho et al., 2020).
We measured the Vela pulsar’s braking index by estimating the frequency second derivative through linear fitting of frequency first derivative measurements just before glitch epochs, yielding a braking index value of 2.94 0.55. As a caution, it may be noted that our estimate assumes almost a full recovery with stable values of the first and second spin derivatives. This value is expected to become more constrained with future observations as this would increase glitch numbers or dataset coverage. It is close to the prediction of the model that invokes the magnetic dipole radiation in a plasma-filled magnetosphere (e.g. Ekşi et al. (2016)). Previous attempts to estimate the braking index for the Vela pulsar by different studies yielded qualitatively different results: a value of 1.4 0.2 (Lyne et al., 2010), 1.7 0.2 (Espinoza et al., 2017) and 2.81 0.12 (Akbal et al., 2017), respectively. The values in each case were different potentially attributed to differences in techniques used to estimate spin parameters and their derivatives and assumptions.
V Conclusion and Future Prospects
The investigation of post-glitch recovery offers valuable insights into neutron star internal structure, which is not fully understood yet, and illuminates the underlying mechanisms driving these sudden spin-up events. Focusing on the Vela pulsar’s last four glitches, spanning from MJD 57651 to 60700 ( 100 months of observations), we employed the vortex creep model to extract key neutron star parameters. Bayesian analysis demonstrated that the recovery of the Vela glitches is well described by a hybrid model combining exponential and linear relaxations. Additionally, Bayesian inference closely predicted the time to the next glitch for all the glitches and suggested that the upcoming glitch is likely to occur around MJD 61377.7 (December 3, 2026), with a 68% credible interval spanning from MJD 61249 to 61506. The vortex creep model provided an excellent fit for the observed recoveries.
Furthermore, we analyzed the vortex residuals (the difference between the vortex creep model predictions and the measured spin-down rate) within the vortex bending model framework, conducting the first study of this kind for the Vela pulsar. Our analysis confirms damped quasi-periodic oscillations and yields precise determinations of the oscillatory parameters. Intriguingly, we also see an unexplained feature in the vortex residuals just preceding the subsequent glitch. These findings provide a foundational understanding of vortex residuals and motivate further enhancements in future both in theory and in similar follow-up investigations. Moreover, the measured parameters can improve the precision of hydrodynamical simulations of superfluid vortices, allowing for more accurate modeling and deeper insights into the neutron star’s dynamics and structural properties.
We explored several correlations and our analysis revealed a positive correlation between the glitch amplitude and the subsequent inter-glitch time, suggesting a potential link between the magnitude of glitches and the time elapsed before the next occurrence. Furthermore, we estimated the pulsar’s braking index as 2.94 0.55. As more glitches are observed, this estimate is likely to become more precise, yielding a more refined understanding of the pulsar’s rotational dynamics and offering deeper insights into its internal structure, magnetospheric conditions and evolutionary history.
Future research includes investigating the post-glitch recovery and internal structure of other notable glitching pulsars, such as the Crab and PSR J1740-3015. To streamline the process, we are developing automated rotational parameter estimation scripts that perform the timing analysis quickly, enabling rapid model investigation and facilitating timely results. Additionally, the measured recovery parameters provide a unique approach to constrain the neutron star equation of state, which will be investigated in a subsequent work in progress. Furthermore, determining the internal and surface temperatures of neutron stars and exploring correlations with glitch observables could yield valuable insights. Lastly, constraining internal magnetic field configurations will also be a key area of investigation, ultimately advancing our understanding of neutron star physics and glitch mechanisms.
Acknowledgements
We acknowledge the support of staff at the Radio Astronomy Centre, Ooty and the upgraded Giant Metrewave Radio Telescope during the observations. The ORT and the uGMRT are operated by the National Centre for Radio Astrophysics. This paper includes archived data obtained through the Parkes Pulsar Data archive on the CSIRO Data Access Portal (http://data.csiro.au). The Parkes radio telescope is part of the Australia Telescope, which is funded by the Commonwealth of Australia for operation as a National Facility managed by the Commonwealth Scientific and Industrial Research Organisation (CSIRO).
We acknowledge the National Supercomputing Mission (NSM) for providing computing resources of ‘PARAM Ganga’ at the Indian Institute of Technology Roorkee, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India.
Funding statement
EG is supported by the National Natural Science Foundation of China (NSFC) programme 11988101 under the foreign talents grant QN2023061004L. BCJ acknowledges the support from Raja Ramanna Chair fellowship of the Department of Atomic Energy, Government of India (RRC – Track I Grant 3/3401 Atomic Energy Research 00 004 Research and Development 27 02 31 1002//2/2023/RRC/R&D-II/13886). BCJ also acknowledges support from the Department of Atomic Energy Government of India, under project number 12-R&D-TFR-5.02-0700. PA acknowledges the support from SERB-DST, Govt. of India, via project code CRG/2022/009359.
Data and Codes Availability
Data and codes used for this research will be provided on reasonable requests.
Appendix A Results of Bayesian Analysis
Here, we present the outcomes of the Bayesian model comparison and the posteriors for the parameters estimated for each post-glitch recovery using the most preferred model. The results of the Bayesian Model comparison are listed in Tables 3 and 4. The posterior distributions are plotted with 68% and 95% credible intervals for all the cases given in Figs 13, 14, 15, 16, 17 and 18.
| Epoch | Model 1a | Model 1b | Model 1c | Model 2 | Model 3a | Model 3b | Model 3c |
| 57734 | 0 | 94237.4 | 74201.3 | 106452 | 135976.6 | 137496.8 | 136553.2 |
| 58515 | 0 | 62619.4 | –370039.7 | –41366.5 | 159957.2 | 152789.7 | 167993.8 |
| 59417 | 0 | 2835.3 | –94.1 | –311756.5 | 223738.7 | 225164.4 | 225196.4 |
| 60429 | 0 | 6590.0 | 6789.0 | –9657.3 | 5796.7 | 6945.6 | 7016.4 |
| Epoch | Hypothesis I | Hypothesis II | Hypothesis III |
| 57734 | 0 | 1350.0 | –3365.7 |
| Peak 1 Shifted | 0 | 730.9 | –3377.2 |
| 59417 | 0 | 4514.4 | –13296.9 |
| Peak 1 Shifted | 0 | –10572.9 | –13302.8 |
| Peak 2 Shifted | 0 | –3800.1 | –13301.3 |