### v2

# Post-glitch Recovery and the Neutron Star Structure: The Vela Pulsar

Abstract
We present a detailed analysis of the Vela pulsar’s rotational behaviour using approximately 100 months of observational data spanning from September 2016 to January 2025, during which four glitches were identified. Here, we demonstrate the post-glitch recovery of these glitches within the framework of the vortex creep model. We further present the investigation of vortex residuals (the discrepancy between observed values and those predicted by the vortex creep model) by interpreting them in the context of the vortex bending model. In addition, we report a positive correlation between the glitch magnitude and the time to the next glitch, applicable only for the large glitch events observed in the Vela pulsar. Furthermore, we estimate the braking index of the Vela pulsar to be 2.94 0.55.
keywords:
radio astronomy; pulsars: Vela; astronomy data analysis; pulsar timing method;doi:
xx.yyyy/pasa.zzzz.tt1 Introduction
Pulsars are highly magnetised, rotating, dense, and compact stars observed to have periodic radiation of pulses with the periodicity tied to their remarkably stable rotation. The slowdown in the rotation of a pulsar is influenced by two types of instabilities: glitches and timing noise (D'alessandro96; Haskell_2015; zhou_erbil_glitch_review). Glitches are abrupt changes in the rotational frequency and spin-down rate of a pulsar, while timing noise is a stochastic wander in the spin parameters of pulsars. The occurrence of glitches, as well as their slow relaxations, which range from a few days to several years, reveals the presence of neutron superfluidity within the inner crust of the neutron star. One of the most widely accepted mechanisms for explaining glitches and their recoveries is the superfluid vortex pinning and creep model (AndersonItoh1975; Alpar1984I; Alpar1984II). This model suggests that neutron superfluidity in the interior of the star leads to a differential rotation, storing a significant amount of angular momentum. When this angular momentum is released, the crust spins up, resulting in a glitch. Pulsar glitches thus provide valuable insights into the superfluid dynamics within the star (zhou_erbil_glitch_review; Liu2024).
The post-glitch recovery is the process in which the pulsar gradually relaxes back towards its original pre-glitch rotational state. According to the vortex creep model, this process is driven by the repinning of superfluid vortices to the crustal nuclei and the re-establishment of the equilibrium lag between the rotational rates of the superfluid interior and the crust. The specific timescale and form of the relaxation, whether exponential or linear, are determined by the distinct vortex dynamics at play. While the standard recovery reflects the transfer of angular momentum from the interior to the surface, deviations from this norm can indicate complex interactions, such as crustal events or non-linear creep regimes (Akbal_2015; Erbil_2023). Ultimately, the post-glitch recovery serves as a probe for the internal dynamics and structure of neutron stars (Erbil_2022).
PSR J0835–4510, B0833–45, the Vela pulsar, has 26 reported glitch events, most of which are large glitches and detected after a regular, quasi-periodic inter-glitch time111http://www.jb.man.ac.uk/pulsar/glitches.html (Espinoza2011; Basu2022). The Vela pulsar is one of the most studied glitching pulsars. The first-ever glitch was recorded in the Vela pulsar (Radhakrishnan_Manchester_1969; Reichley_Downs_1969). Moreover, the first-ever glitch detection caught in the act was also realised for the Vela pulsar (Palfreyman2018_Vela). The Vela pulsar is considered one of the most popular glitching pulsars for several reasons: it is a bright, young source that displays glitches at regular intervals. The recoveries of glitches also demonstrate strikingly similar behaviour. A measure of a pulsar’s slowdown over time is described by the braking index (). It is defined by a power-law relation between the pulsar’s spin-down rate and its rotational angular frequency (), given by . The braking index provides insights into the pulsar’s rotational evolution, magnetic fields, and emission mechanisms.
Here, we present the observational behaviour of the Vela pulsar from September 2016 to January 2025; during this period, four large glitches were observed in our monitoring campaign (Grover2024; Basu2020). Theoretical inferences from the post-glitch recovery phase are also presented. We have used the recovery behaviour of four glitches in the Vela pulsar to estimate the following parameters: (a) the fractional moments of inertia (FMI) of various layers participating in a glitch, (b) the re-coupling timescale of the crustal superfluid, (c) theoretical prediction for the time to the next glitch, (d) braking index of the Vela pulsar. The paper’s outline is as follows: a brief description of the vortex creep and bending models is given in Section 2. We describe the pulsar observations, the data reduction, and the analysis processes in section 3. In Section 4, we present the results. The conclusions and future scope are given in Section 5.
2 The Vortex Creep and Bending Models
The vortex creep model provides a microscopic framework for understanding the macroscopic behaviour of glitches and their subsequent recovery. In the neutron star crust, vortex lines coexist with and energetically pin to the crystal lattice. Crustal inhomogeneities such as dislocations, cracks, and impurities produced by quakes create a non-uniform vortex distribution known as vortex traps (Cheng1988). Glitches may arise due to collective vortex unpinning events induced in these vortex traps when the angular velocity lag between the superfluid and the crustal normal matter exceeds a threshold value (AndersonItoh1975; Alpar1984I). During the unpinning events, angular momentum is rapidly transferred from the superfluid to the crust. The superfluid regions involved in the unpinning decouple from the external braking torque, resulting in a step increase in the spin-down rate.
After a glitch, the steady-state lag decreases, temporarily suppressing creep in some superfluid regions. The recovery process involves restoring the creep until steady-state conditions are re-established by the external braking torque. Shortly after the glitch, excess angular momentum is transferred from the crustal superfluid to the normal matter in the crust, producing a transient peak in the crustal rotation rate. This perturbation is subsequently communicated to the core superfluid through electron scattering off magnetised vortex lines (Alpar1984II), causing the initial peak to relax on the crust–core coupling timescale toward a new equilibrium state. A pronounced peak in the rotation rate of the normal component is expected when the coupling between the crustal superfluid and the normal matter is stronger than that with the core superfluid (Pizzochero2020). On longer timescales, the core superfluid recouples, leaving only the crustal superfluid effectively decoupled from the external braking torque.
In some glitches of the Crab pulsar, delayed spin-ups are observed (Wong2001; Shaw2018; Basu2020; Ge_2020). Within the vortex creep framework, these features are attributed to vortex unpinning accompanied by inward vortex motion driven by fractured crustal platelets (Akbal_2015; Erbil_2019). Additionally, damped, sinusoidal-like oscillations in the spin-down rate can arise from vortex-line bending in the non-linear creep relaxation regime, excited by the glitch and potentially influenced by preceding events (Erbil_2023). It is important to note that the vortex creep model is not the unique explanation; several equally viable models successfully explain glitch-triggering mechanisms, delayed spin-ups, and post-glitch relaxation. For a comprehensive discussion, see Haskell_2015; Antonopoulou2022; Antonelli2022; zhou_erbil_glitch_review.
The post-glitch spin-down rate behaviour mirrors the responses of the vortex creep regions in both linear and non-linear regimes to a glitch, with additional involvement of induced inward vortex motion in delayed spin-up events. In certain regions of the superfluid in the inner crust and the outer core, vortex creep exhibits a linear response to glitch-induced changes, leading to exponential relaxation (Alpar1989; Flanagan1990; Erbil_2017aug) as:
where , is the spin frequency, is the spin frequency just before the glitch, is the FMI of the of ith layer that participated in a glitch, and is the corresponding exponential recovery timescale.
The overall linear relaxation, obtained by integrating the contributions of several neighbouring single non-linear regime regions (Alpar1984I; Erbil_2022), with an assumption of linearly decreasing superfluid angular velocity during a glitch (Alpar1996) and is given as:
where is the FMI of the non-linear crustal superfluid region, is the offset time, and is the non-linear creep relaxation time. The non-linear creep response gives an estimate for the time to the next glitch. The glitch-induced changes in the spin-down rate fully relax after a waiting time (Erbil_2022),
where is the observed glitch magnitude and represents the decrement in the superfluid rotational frequency. The post-glitch recovery completes once the relaxation of the non-linear creep regions is finished. This gives us a theoretical prediction for the time to the next glitch in terms of the vortex creep model, which lies between the range to .
A typical post-glitch recovery involves an exponential and a linear relaxation described by equations 1 and 2, respectively. This is tested by using different alternative models as described below:
Model 1: Only exponential recovery given by equation 1.
Model 1a: Exponential recovery with only 1 exponential component.
Model 1b: Exponential recovery with 2 exponential components.
Model 1c: Exponential recovery with 3 exponential components.
Model 2: Only linear relaxation described by equation 2.
Model 3: Exponential and linear recoveries, the sum of equations 1 and 2.
Model 3a: The sum of linear recovery and exponential recovery with only 1 exponential component.
Model 3b: The sum of linear recovery and exponential recovery with 2 exponential components.
Model 3c: The sum of linear recovery and exponential recovery with 3 exponential components.
The most preferred model and corresponding parameters among the various alternatives were evaluated using Bayesian analysis of our post-glitch recovery measurements of the Vela pulsar. Further details of this analysis are presented in the next section.
Apart from long-term post-glitch recovery, we see evidence for sinusoidal-like spin-down rate oscillations over and above the recovery predicted by the vortex creep model as described in Section 4. Such oscillations can be explained by vortex bending (Erbil_2023) and appear to be a common feature following glitches, as also observed from PSRs J1522–5735 (zhou24), J0742–2822 (zubieta25), J1509–5850, and J1718–3825 (liu25). To understand this better, we used three alternative hypotheses as described below
Hypothesis I – An undamped model consisting of only sinusoidal term, .
Hypothesis II – Damped model consisting of a sinusoidal term and a damping factor, .
Hypothesis III – An oscillatory model as described in Erbil_2023 consisting of .
where is the amplitude, is the angular frequency, is the phase, and is the decay time for the oscillation. The next section outlines the observations, data reduction, and timing used to derive the spin frequency and spin-down rate. It also describes the Bayesian analysis used to calculate the evidence for the competing hypotheses, thereby identifying the most plausible model and estimating its corresponding parameters.
3 Observations and Analysis
3.1 Sample and Observational Facilities
We use two large radio telescopes to monitor pulsars: the upgraded Giant Metrewave Radio Telescope (uGMRT)(ugmrt2017) and the Ooty Radio Telescope (ORT) (swarup1971ort). Observations of PSR B083345 were part of a regular glitch monitoring program, where we regularly monitor 24 pulsars (i) to detect glitches, (ii) model timing noise, and (iii) model post-glitch recovery. This work presents a comprehensive analysis of the Vela pulsar’s rotational evolution, spanning from September 2016 to January 2025, as part of our ongoing monitoring campaign (Grover2024; Basu2020). Our study primarily uses observations from the ORT, supplemented by uGMRT and archival ATNF data to fill temporal gaps and facilitate specialised investigations.
The ORT (swarup1971ort) is a 530m 30m parabolic cylindrical reflector. It is located on a hill with the same slope as the geographic latitude (11°), allowing it to track sources for about 8 hours and covers a declination range of -60° to +60°. Pulsar observations at the ORT are carried out at the central observation frequency of 326.5 MHz using the Pulsar Ooty Radio Telescope’s New Digital Efficient Receiver, PONDER (Naidu_2015), which produces real-time coherent dedispersed time series data. The typical cadence of the observations at the ORT is 1–14 days.
The GMRT is an interferometer with 30 parabolic antennas, with a diameter of 45 meters each (Swarup1991GMRT; Ananthakrishnan1995GMRT). The uGMRT (ugmrt2017) covers a declination range of -53°to +90°and supports observations from 150 MHz to 1460 MHz in four bands. The Vela pulsar is observed in uGMRT Band 3 (250–500 MHz) and Band 4 (550–950 MHz). The central square antennas, together with a subset of the nearest arm antennas, were used in a phased array to synthesise a large single dish for observations. The outer arm antennas were not included in the phased array due to significant drift in the phase of the voltage outputs, caused by ionospheric variations, in these antennas. A digital filterbank, implemented in the GMRT correlator, was used to record either 1024 or 2048 channels across the bandpass to obtain frequency-resolved data with a typical sampling time of 655 s. The cadence of the observations is 15-30 days. We primarily use frequency-resolved uGMRT data to examine the impact of dispersion measure (DM) offsets on the spin-down rate. Additionally, uGMRT data filled a temporal gap around epoch 60300, providing crucial estimates not covered by ORT due to maintenance downtime.
The Commonwealth Scientific and Industrial Research Organisation (CSIRO) Data Access Portal is used to access the Parkes pulsar archival observations222https://data.csiro.au/domain/atnf (Hobbs2011) of the Vela pulsar. These data were mainly used to fill temporal gaps in glitch number 2 (MJD 58515), caused by significant observational gaps at ORT. We directly downloaded and used the reduced files (or PSRFITS: PSRCHIVE_and_PSRFITS2004) and performed subsequent analysis, as described in the next subsection.
3.2 Data Reduction and Timing Analysis
We used the DSPSR software (DSPSR) and the PINTA pipeline (pinta) to reduce the raw data from the Vela pulsar observed at the ORT and the uGMRT, respectively, into PSRFITS files, using the ephemeris created in our initial observations. The initial analysis, including the generation of a noise-free template and Time of Arrivals (ToAs), was performed using PSRCHIVE (PSRCHIVE_and_PSRFITS2004; PSRCHIVE2012). The ToAs were obtained through the cross-correlation of the template profile with all other observed profiles using the ‘pat’ command, using the Fourier Phase Gradient (PGS) method (Taylor1992). The precision timing analysis of the pulsar was done with the help of the pulsar timing software TEMPO2 (Hobbs2006Tempo2; Edwards2006Tempo2). The DM offsets were estimated using the DMcalc (Krishnakumar2021) script from the frequency-resolved data from uGMRT. The rotational evolution of a pulsar including a glitch can be expressed as a Taylor series expansion (Erbil_2022),
where is the spin frequency, and and are the first and second derivatives of the frequency, respectively, at reference epoch . and denote the changes in spin frequency and its first derivative at the glitch epoch , respectively. The parameters and correspond to the amplitude and characteristic decay timescale of the exponential relaxation component following the glitch.
The spin evolution of the Vela pulsar from September 2016 to January 2025, obtained after precision timing, is given in Fig. 1. Each measurement is made with the local timing of several epochs, typically within a 50-200 day data window, contingent on data availability. Exceptions include the first and last few points adjacent to glitches. The resulting measurements are reflected in the precision of error bars, which are smaller than the symbol used in the figure. By choosing multi-epoch windows, we achieve reliable rotational estimates with minimised error bars, providing a robust foundation for studying vortex residuals. This approach effectively averages out temporal effects caused by physical processes, such as ionised interstellar medium scattering-induced broadening, thereby validating the accuracy of measurements aligned with the star’s rotation and internal dynamics. These recovery measurements are investigated within the framework of the vortex creep and bending models as described in Section 2 and the results are discussed in Section 4.
3.3 Bayesian Analysis
For each recovery, we used Bayesian statistics for model comparison and parameter estimation. Bayesian statistics use Bayes’ theorem to enhance the probability of a hypothesis based on new evidence or data. We used DYNESTY (a dynamic nested sampling package for estimating Bayesian posteriors and evidences; (DYNESTY2020)) to compute the Bayesian evidence and perform model selection, and EMCEE (EMCEE2013) to obtain posterior samples for parameter estimation of the preferred model. For model comparison, the standard Jeffreys1961 scale is used, which suggests the preferred model based on the strength of evidence (Trotta2008). Bayesian inference is also used to predict the epoch of the next glitch using data from previous glitches.
The priors are selected based on the initial knowledge of each parameter, as informed by the literature (Erbil_2022; Flanagan1990; Alpar1989). This knowledge provides reasonable initial ranges that can be further refined or broadened based on the results. For instance, parameters such as the fractional moment of inertia ( or ) is generally positive and below unity. The recoupling timescale in exponential regions () depends on the number of exponential components (three-, two-, or one-component models). Typically, for the Vela pulsar, it has been observed that the first component lies within 0.1–10 days, with the second and third components approximately 10 and 100 times longer, respectively.
For the linear relaxation region, the prior for recoupling timescale () is around 10-300 days, and the offset time () is 300 to 1200, with the initial guess often chosen close to the observed time to the next glitch. For the Vortex bending studies, the prior on amplitude, phase, and decay constant were set as 0–1, 0–2, and 0–2000 days, respectively. All the priors were considered as uniform priors and listed in Table 1. While these priors adequately describe many recoveries, particularly those observed in the Vela pulsar, their applicability may vary depending on the nature of the recovery. As each recovery exhibits unique characteristics, a flexible and adaptive approach to defining priors is essential. We model the data assuming independent Gaussian measurement uncertainties. The corresponding log-likelihood function is given by
where denotes the observed data, is the model prediction for parameters , and represents the corresponding measurement uncertainty.
| Parameters | Values | Parameters | Values |
| 0–1 | 0–1 | ||
| 0–1 | 10–300 days | ||
| 0–1 | 300–1200 days | ||
| 0.1–10 days | 0–1 | ||
| 5–100 days | 0–2 | ||
| 10–400 days | 0–2000 days |
The post-glitch recoveries are studied within the framework of the vortex creep model and the vortex bending model (Erbil_2023), discussed in the previous section. A quick picture of the analysis process is as follows: We first model the frequency derivative measurements from the data using the vortex creep model. Then, we estimate the residuals obtained by subtracting the predicted values from the vortex creep model from the measured values. These residuals were investigated within the framework of the vortex bending model. The generalised Lomb-Scargle periodogram technique (Lomb1976; Scargle1982; VanderPlas2018) was employed to detect periodicities in the residuals. We used the Astropy (Astropy2013) implementation of the generalised Lomb-Scargle periodogram. The Peaks with false alarm probabilities (FAP) below 0.1% or Z-score above 3 were considered significant for vortex bending studies. FAP estimation was performed using the Baluev2008 method. Alternative methods, including approximations based on effective independent frequencies and bootstrap resampling, yielded similar results. The significance of each detected peak is quantified using the Lomb–Scargle false-alarm probability (FAP). For interpretive purposes, we also report the equivalent z-score values (Cowan2011; Ganguly_Desai_2017). The results of our analysis are presented in the next section.
4 Results and Discussions
Our pulsar monitoring program with the ORT and the uGMRT started in 2015. The previous results were reported in Basu2020 and Grover2024. In this work, we present an analysis of a new Vela glitch at MJD 60429.9. Further, we demonstrate the spin-down rate evolution for the Vela pulsar since the beginning of the monitoring program. Our primary focus is to provide the theoretical interpretation of glitches using the post-glitch recovery phase and to obtain measurements for the re-coupling timescales and fractional moment of inertia for various layers that participate in a glitch. We have estimated the theoretical time to the next glitch through the vortex creep model and used Bayesian inference to predict the time for the next glitch. All the predictions match well with the observed inter-glitch times.
4.1 Glitches
Here, we summarise the last three glitches in the Vela pulsar: MJD 57734.4, 58517, and 59417.6, detected in our monitoring program (Grover2024; Basu2020). The recoveries for each of these glitches are studied within the framework of the vortex creep model to investigate the structure of the Vela pulsar. Furthermore, we report the detection and analysis of the recent large glitch in the Vela pulsar in 2024. The characteristics and parameters of glitches are listed in Table 2.
| Number | Epoch (MJD) | Pre-glitch (Hz) | Pre-glitch (Hz s-1) | () | Reference | |
| G1 | 57734.4(2) | 11.186433252(6) | -1.556383(8) | 1433.2(9) | 5.595(9) | Basu2020 |
| G2 | 58517(7) | 11.185397242(3) | –1.555659(4) | 2471(6) | 6(2) | Grover2024 |
| G3 | 59417.6(1) | 11.184208478(5) | –1.5550(8) | 1235(5) | 8.0(7) | Grover2024 |
| G4 | 60429.9(1) | 11.182859362(3) | –1.5530(8) | 2396(2) | 23(1) | This work |
We present the parameters of the Vela 2024 glitch, which occurred around MJD 60429.9. This glitch has been reported by several groups (Zubieta2024ATel16608_Vela; Campbell-Wilson2024ATel16610_Vela; Grover2024ATel16611_Vela; Palfreyman2024ATel16615_Vela; Wang2024ATel16619_Vela; Zubieta2025_Vela). Here, we present updates in our preliminary results reported in Grover2024ATel16611_Vela. The glitch is estimated to occur at MJD 60429.9(1), with the fractional increase in the spin frequency, and its time derivative is calculated as and , respectively. The results of the glitch epoch and spin parameters agree with the reported estimates of other monitoring programs. A more precise measurement of glitch epoch as 60429.86962(4) is also presented in Palfreyman2024ATel16615_Vela with higher cadence observations. The timing residuals, the evolution of the spin frequency, and the spin-down rate are given in Fig. 2. The post-glitch recovery for this glitch is also investigated within the framework of the vortex creep model in the next subsection.
4.2 Post-Glitch Recoveries
In this section, we present the post-glitch recovery behaviour of the Vela pulsar. The spin-down rate has been analysed from MJD 57651 to 60700. During this period, the Vela pulsar has displayed four glitches. The post-glitch recovery of the Vela pulsar is regular, i.e., it follows similar behaviour. The recovery for this pulsar usually consists of exponential and linear relaxations well described by Model 3 of Section 2. We have investigated all post-glitch recoveries for this pulsar within the framework of the vortex creep model for the glitches listed in Table 2, and the resulting fit parameters obtained using the vortex creep model are listed in Table 3. Now, we concisely describe the post-glitch relaxation(s) for individual cases.
| Epoch | |||||||||||
| (MJD) | () | (days) | () | (days) | () | (days) | () | (days) | (days) | (days) | (days) |
| 57734.4(2) | 783 | [787, 1083] | |||||||||
| 58517(7) | — | — | 902 | [752, 1033] | |||||||
| 59417.6(1) | 1012 | [855, 1123] | |||||||||
| 60429.9(1) | — | [820, 1076] |
4.2.1 Glitch No. 1 – MJD 57734.4(2)
G1, or glitch at MJD 57734.4(2) (Sarkissian2017; Palfreyman2018_Vela; Ashton2019; Basu2020; Erbil_2020; Montoli2020; Dunn2025), was detected with a fractional change in the rotational frequency of 1433(1), and a fractional change in the spin-down rate as 5.59(1). We present the post-glitch relaxation modelling for this glitch in Fig. 3. For this recovery, the Bayesian model comparison prefers the presence of three-exponential + linear recovery (Model 3c) over all other models. The estimated median values for parameters are: the fractional moment of inertia for the first, second, and third exponential components are , and respectively, with the decay timescales of , and days, respectively. The median fractional moment of inertia for the non-linear creep region is with a recoupling timescale of 62 days. The offset time is measured as days, close to the observed inter-glitch time of 783 days. The observed time is consistent with the vortex creep model’s predicted range and closely matches the Bayesian prediction, which estimates a 68% credible interval of 787 to 1083 days. The fit parameters, along with their 95% credible intervals, are listed in Table 3.
The bottom part of Fig. 3 shows the residuals obtained after subtracting the vortex creep model predictions from the observed data and is in units of Hz s-1. These vortex residuals, spanning from 200 days onward, were used for vortex bending studies, excluding the initial 200 days due to the significant impact of exponential recoveries. The generalised Lomb–Scargle periodogram (Fig. 4) reveals a prominent peak corresponding to a period of 315(1) days using the least-squares interpretation of the periodogram (VanderPlas2018). This signal exhibits strong significance, with a false alarm probability (FAP) of (Z-score: 5.0). A secondary peak at approximately 165 days has a FAP of 0.14, while all remaining peaks yield FAP values of 1, indicating low significance.
Fixing the period at 315(1) days, we investigated the vortex bending hypothesis and found that the damping model, comprising a sinusoidal term and a damping factor (Hypothesis II), was strongly favoured over Hypotheses I and III. Additionally, we examined whether the observed peak resulted from shocks caused by the current glitch or the previous one, concluding that the oscillations decisively originated from the current glitch. The comparison of vortex residuals and vortex bending model is shown in Fig. 5 and the estimated parameters for the damped model (Hypothesis II) are: Amplitude = , Phase = , and days. Note that vortex residuals are in scale of Hz s-1, so the effective amplitude is .
The vortex residuals fit with the vortex bending model well for around 200-400 days post-glitch duration, but deviate thereafter. Including a second peak improves the fit, but the generalised Lomb-Scargle periodogram indicates a FAP of 14% for this peak using the Baluev method (with a minimum FAP of 5.3% obtained through approximations based on effective independent frequencies). Given that the FAP of the second peak is greater than 0.1%, we exclude the second peak from our analysis. The other peaks with a high FAP are likely to originate from preceding glitches, and their effect on the present glitch recovery is insignificant. The post-glitch recovery modeling for this pulsar is also presented by Erbil_2022 using the Fermi-LAT observations. However, they only detected one exponential component. We are reporting three exponential recoveries made possible by much higher cadence observations, which allow us to investigate the vortex residuals within the framework of the vortex bending model.
4.2.2 Glitch No. 2 – MJD 58517(7)
The glitch at MJD 58517(7), G2 in the Vela pulsar, (Sarkissian2019ATel12466; Kerr2019ATel12481; LopezArmengol2019ATel12482; lower2020_utmost2; Erbil_2022; Grover2024; Dunn2025), was observed with a fractional change in the rotational frequency of 2471(6), and a fractional change in the spin-down rate as 6(2). Our analysis focuses on modelling the post-glitch relaxation behaviour, given in Fig. 6. Our ORT and uGMRT observations weren’t conducted during the initial exponential relaxation phase of this glitch, and we detected only non-linear creep recovery. To proceed with our post-glitch analysis, we supplemented our dataset with publicly available ATNF pulsar data to fill temporal gaps. Additionally, we used estimates derived from analysis using data closer to the glitch from lower2020_utmost2, specifically the glitch epoch (58515.6) and the fractional change in rotational frequency and its derivative (2501 and 8.7, respectively), for our recovery studies.
The Bayesian model comparison reveals that the model with two exponential terms and a linear recovery term (Model 3b) provides the best description of this recovery. The measured median values for the recovery parameters are as follows: the fractional moment of inertia for the second and third exponential components are and respectively, with the decay timescales of and days respectively. The fractional moment of inertia for the non-linear creep region is with a recoupling timescale of days. The offset time is estimated as days. The observed inter-glitch time of 902 days is consistent with both the predicted range of the vortex creep model and Bayesian statistics. The median value predicted by Bayesian statistics is 893 days, very close to the observed time. These parameters are also summarised in Table 3. Unlike other glitches that were observed with very high cadence using the ORT, this glitch was not monitored with such precision, and therefore, the vortex-bending oscillations are not well observed and studied in this work. Our post-glitch recovery modelling results differ from those of Erbil_2022, who reported a single exponential component using Fermi-LAT observations. Our analysis benefits from a higher cadence and reveals two well-constrained, distinct exponential recoveries.
4.2.3 Glitch No. 3 – MJD 59417.6(1)
G3, or the glitch at MJD 59417.6(1) (Sosa-Fiscella2021ATel14806; Dunn2021ATel14807; Olney2021ATel14808; Singha2021ATel14812; Zubieta_2023; Grover2024; Dunn2025) was observed to have a fractional spin up in frequency and a fractional spin up in frequency-time derivative . We present the post-glitch relaxation behaviour for this glitch in Fig. 7. In addition to the ORT, we leverage observed data from the uGMRT as well, with two motivations: (i) to confirm that the residuals stem from the neutron star’s interior, rather than the ionised interstellar medium by using DM offsets in the uGMRT analysis and (ii) to fill the temporal gap between 850 to 1000 days subsequent to the glitch date.
High-cadence ORT observations of this glitch enable us to model three exponential components present in the recovery; Model 3c emerged as the top choice in Bayesian model comparisons, outperforming all other models. The calculated median values for the parameters are as follows: the fractional moment of inertia for the first, second and third exponential components are and and respectively, with the decay timescales of , and days, respectively. The fractional moment of inertia for the non-linear creep region is with a recoupling timescale of days. The offset time is estimated as days, which deviates significantly from the observed inter-glitch time primarily due to the large value of (equivalent to ). However, the vortex creep model predicted time for the next glitch lies between to , which aligns well with the observed time. The median Bayesian prediction of 989 days is near the observed interval of 1012 days, falling well within the 68% credible interval. These measured parameters are listed in Table 3.
The relaxation behaviour of this glitch has never been investigated before. The lower panel of Fig. 7 shows the vortex residuals, obtained after subtracting the vortex creep model predictions from the observed data, in scale of Hz s-1. We analysed residuals between 200 and 700 days for vortex bending studies, excluding the initial 200 days due to dominant exponential recovery effects, and data beyond 700 days requires further investigation in existing theory.
The Lomb–Scargle periodogram analysis of the vortex residuals (Fig. 8) reveals several peaks, but only the first two are statistically significant. The strongest peak yields a FAP of 4.6 10-13 and a Z-score of 7.2, while the latter peak has a FAP of 2.46 10-5 and a Z-score of 4.1. The third peak, with a period of around 81 days, has a FAP of 0.88 and is therefore not significant; all remaining peaks have FAP values consistent with unity. The corresponding periods of the two significant peaks are approximately 344.0(7) and 153.5(7) days, estimated using the least-squares interpretation of the periodogram (VanderPlas2018).
These two peaks are used to investigate the vortex residuals within the framework of the vortex bending hypothesis, presented in Fig. 9. We observed that Hypothesis II was strongly favoured over Hypotheses I and III. Additionally, we examined whether the observed peak resulted from shocks caused by the current glitch or the previous one, concluding that the oscillations decisively originated from the current glitch. The estimated parameters for the damped model (Hypothesis II) are: (i) Amplitudes, A1 = , and A2 =, (ii) Phase, and , (iii) and days. Note that vortex residuals are in scale of Hz s-1, so the effective amplitudes are , and , respectively.
4.2.4 Glitch No. 4 – MJD 60429.9(1)
We present the post-glitch relaxation for the recent glitch at MJD 60429.9(2), G4 (Zubieta2024ATel16608_Vela; Campbell-Wilson2024ATel16610_Vela; Grover2024ATel16611_Vela; Palfreyman2024ATel16615_Vela; Wang2024ATel16619_Vela; Zubieta2025_Vela). The glitch is illustrated in Fig. 2, and the comparison of the observed data with the vortex creep model fitting is presented in Fig. 10.
The Bayesian model comparisons selected Model 3c as the best model over all other models. The calculated median values for the parameters are as follows: the fractional moment of inertia for the first, second and third exponential components are and and respectively, with the decay timescales of , and days, respectively. The median fractional moment of inertia for the non-linear creep region is with a recoupling timescale of days. The median value for the offset time is days. These parameters, along with their 95% credible intervals, are listed in Table 3. The parameters for the linear creep regions (or exponential recovery) are well-constrained. However, those dependent on the non-linear creep region will likely be refined in the future as more data becomes available. The Bayesian analysis forecasts a median time for the next glitch at MJD 61377.7, with a 68% credible interval spanning from MJD 61249 to 61506.
4.3 Correlations and Braking Index
To date, 26 glitches have been reported in the Vela pulsar333http://www.jb.man.ac.uk/pulsar/glitches.html (Espinoza2011; Basu2022), with 21 of them having large magnitudes (¿ 10-7). Fig. 11 presents these 26 events. A pattern is observed among the large glitches, where their magnitudes exhibit an alternating trend: a glitch with a lower magnitude is typically followed by one with a higher magnitude, and vice versa. This alternating behaviour is particularly evident in the last 10 glitches (i.e., after MJD 50000), as highlighted by the red boundary in Fig. 11.
We used the reported data of 26 glitch events to investigate potential correlations between the glitch magnitude and the inter-glitch time. A significant correlation was observed between the glitch magnitude and subsequent inter-glitch time specifically for events with large magnitudes, shown in Fig. 12. The correlation coefficients were: Pearson’s = 0.6470 (P–value = 0.0020), Spearman’s = 0.5852 (P–value = 0.0067), and Kendall’s = 0.4274 (P–value = 0.0086), indicating moderate to strong positive correlation. Such a correlation has only been found for PSR J0537–6910 (ho20).
This correlation between glitch magnitude and the subsequent interglitch time does not persist when small glitches are included, as illustrated in Fig. 13. The reality of reported small glitch events remains uncertain; some may be consistent with timing noise (Espinoza2011; Grover2024), although it is unlikely that all small glitches are timing noise, especially those accompanied by exponential recoveries (Espinoza2021). Therefore, the interpretation of this correlation should be approached with extreme caution. If the features and correlations shown in Figs. 11 and 12 persist in including all real glitches; this would have significant implications for glitch forecasting models, such as those proposed by Antonelli2022.
We measured the Vela pulsar’s braking index by estimating the frequency second derivative through linear fitting of frequency first derivative measurements just before glitch epochs, yielding a braking index value of 2.94 0.55. As a caution, it may be noted that our estimate assumes almost a full recovery with stable values of the first and second spin derivatives. This value is expected to become more constrained with future observations as this would increase glitch numbers or dataset coverage. It is close to the prediction of the model that invokes the magnetic dipole radiation in a plasma-filled magnetosphere (e.g. eksi16). Previous attempts to estimate the braking index for the Vela pulsar by different studies yielded qualitatively different results: a value of 1.4 0.2 (Lyne1996), 1.7 0.2 (Espinoza2017) and 2.81 0.12 (Akbal2017), respectively. The values in each case were different, potentially attributed to differences in techniques used to estimate spin parameters and their derivatives and assumptions.
5 Conclusion and Future Prospects
Focusing on the Vela pulsar’s last four glitches, spanning from MJD 57651 to 60700 ( 100 months of observations), we employed the vortex creep model to infer model parameters relevant to the post-glitch evolution. Bayesian analysis demonstrated that the recovery of the Vela glitches is well described by a hybrid model combining exponential and linear relaxations. Additionally, Bayesian inference closely predicted the observed time to the next glitch for all the glitches and suggested that the upcoming glitch is likely to occur around MJD 61377.7 (December 3, 2026), with a 68% credible interval spanning from MJD 61249 to 61506.
The vortex creep model provides a good description of the observed post-glitch recoveries. We further analysed the vortex residuals, defined as the difference between the model predictions and the measured spin-down rate, within the vortex bending framework. This analysis identifies damped quasi-periodic oscillations and allows estimation of the associated oscillatory parameters. We also observe an additional feature in the vortex residuals preceding the subsequent glitch, the origin of which remains unclear. These results motivate further theoretical and observational follow-up studies.
We explored several correlations and found a positive relationship between glitch magnitude and the subsequent inter-glitch interval, but only for large glitches. This suggests a possible connection between the magnitude of a glitch and the time until the next occurrence. However, this correlation does not hold when small glitches are taken into account. Consequently, a more comprehensive and robust study, with careful verification of small glitch events, is required. Furthermore, we estimated the pulsar’s braking index as 2.94 0.55.
Future research includes investigating the post-glitch recovery and internal structure of other notable glitching pulsars, such as the Crab and PSR J1740–3015. To streamline the process, we are developing automated rotational parameter estimation scripts that perform the timing analysis quickly, enabling rapid model investigation and facilitating timely results.
Acknowledgements
We acknowledge the support of staff at the Radio Astronomy Centre, Ooty and the upgraded Giant Metrewave Radio Telescope during the observations. The ORT and the uGMRT are operated by the National Centre for Radio Astrophysics. This paper includes archived data obtained through the Parkes Pulsar Data archive on the CSIRO Data Access Portal (http://data.csiro.au). The Parkes radio telescope is part of the Australia Telescope, which is funded by the Commonwealth of Australia for operation as a National Facility managed by the Commonwealth Scientific and Industrial Research Organisation (CSIRO).
We acknowledge the National Supercomputing Mission (NSM) for providing computing resources of ‘PARAM Ganga’ at the Indian Institute of Technology Roorkee, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India.
Funding statement
EG is supported by the Doctor Foundation of Qingdao Binhai University (No. BJZA2025025). BCJ acknowledges the support from Raja Ramanna Chair fellowship of the Department of Atomic Energy, Government of India (RRC – Track I Grant 3/3401 Atomic Energy Research 00 004 Research and Development 27 02 31 1002//2/2023/RRC/R&D-II/13886). BCJ also acknowledges support from the Department of Atomic Energy Government of India, under project number 12-R&D-TFR-5.02-0700. PA acknowledges the support from SERB-DST, Govt. of India, via project code CRG/2022/009359.
Data and Codes Availability
Data and codes used for this research will be provided on reasonable request.
Appendix A Results of Bayesian Analysis
Here, we present the outcomes of the Bayesian model comparison and the posteriors for the parameters estimated for each post-glitch recovery using the most preferred model. The results of the Bayesian Model comparison are listed in Tables 4 and 5. The posterior distributions are plotted with 68% and 95% credible intervals for all the cases given in Figs 14, 15, 16, 17, 18 and 19.
| Epoch | Model 1a | Model 1b | Model 1c | Model 2 | Model 3a | Model 3b | Model 3c |
| 57734 | 0.0 | –14.6 | –16.6 | 126215.8 | 136246.5 | 137750.8 | 137796.8 |
| 58515 | 0.0 | 28647.6 | 28644.6 | 115435.8 | 301825.3 | 309848.9 | 309846.8 |
| 59417 | 0.0 | –13.9 | 71.3 | 221618.3 | 224392.2 | 224606.3 | 224627.1 |
| 60429 | 0.0 | 4451.1 | 4456.2 | 2458.6 | 14541.7 | 15608.3 | 15626.8 |
| Epoch | Hypothesis I | Hypothesis II | Hypothesis III |
| 57734 | 0.0 | 1283.0 | –3441.4 |
| Peak 1 Shifted | –0.1 | 886.1 | –3457.2 |
| 59417 | 0.0 | 3432.4 | –14222.3 |
| Peak 1 Shifted | –4.9 | 618.1 | –23907.7 |
| Peak 2 Shifted | –37.0 | 1581.4 | –14224.4 |